# Wasserstein Proximal Coordinate Gradient Algorithms

### v1

@@ 2.2 Convex functionals on Wasserstein space
only requires the $\lambda$-strong convexity to guarantee its exponential convergence rate [6]. Note that a proximal gradient algorithm is also called an implicit gradient descent algorithm, since the first-order optimality condition (FOC) for (8) reads $x^{k+1}_{\text{prox}}=x^{k}_{\text{prox}}-\tau\nabla f(x^{k+1}_{\text{prox}})$. The theoretical advantage of the proximal scheme comes at the price of the additional computational cost of numerically solving this FOC equation.
@@
where $\{\mu_{t}:0\leq t\leq 1\}$ is the constant speed geodesic curve connecting $\mu_{0}$ and $\mu_{1}$. When $\lambda=0$, we simply say that $\mathcal{F}$ is geodesically convex. It can be shown that both functionals $\mathcal{W}_{j}(\rho_{j})$ and $\mathcal{H}_{j}(\rho_{j})$ in our problem formulation (4) are geodesically convex when their defining functions $W_{j}$ and $h_{i}$ are convex [2, Chapter 9.3]. However, different from the Euclidean case where all convex functions are continuous, geodesically convex functionals may not be continuous with respect to the $\operatorname{W}_{2}$ distance.
Example 2.1 (Geodesically convex functional may not be continuous).
@@
with initial distribution $\rho^{0}$ and step size $\tau>0$. This is analogous to the proximal gradient algorithm (8), i.e., the backward time-discretization of a gradient flow on the Euclidean space. The convergence of Wasserstein proximal gradient scheme for the KL divergence functional $\mathcal{F}_{\operatorname{KL}}(\rho)=\int V\,{\rm d}\rho+\int\rho\log\rho$ is well studied in literature. As on the Euclidean space, piecewise interpolation $\rho_{t}=\rho^{\lfloor\frac{t}{\tau}\rfloor}$ converges to a continuous dynamics on Wasserstein space, i.e., the solution of Fokker–Planck equation $\partial_{t}\rho_{t}=\nabla\cdot(\rho_{t}\nabla(V+\log\rho_{t}))$ [42]. Since the Fokker–Planck equation is also the density evolution equation of Langevin dynamics [27]
@@
one can also use the Euler–Maruyama method to numerically approximate the Wasserstein proximal gradient scheme for the KL using particle approximation. It is shown in the literature that when $V$ is strongly convex, the solution $\{\rho_{t}:\,t\geq 0\}$ of the Fokker–Planck equation converges to the unique global minimizer $\rho^{\ast}\,\propto\,e^{-V}$ of $\mathcal{F}_{\rm{KL}}$ [48]. Corollary 2.8 in [53] further shows that the time-discretization of the continuous time dynamic via Wasserstein proximal gradient scheme $\{\rho^{k}:k\in\mathbb{Z}_{+}\}$ is unbiased, and also converges to $\rho^{\ast}\propto e^{-V}$ exponentially fast when $V$ is strongly convex.
@@ 2.3 First variation and Wasserstein gradient
for any perturbation $\chi=\tilde{\mu}-\mu$ such that $\tilde{\mu}\in\mathscr{P}_{2}^{r}(\mathcal{X})$ and $\int\,{\rm d}\chi=0$. Clearly, $\frac{\delta\mathcal{F}}{\delta\rho}(\mu)$ is uniquely defined up to additive constant. This is analogous to the gradient of $\mathcal{F}$ in the $L^{2}(\mathcal{X})$ sense. If $\mathcal{F}$ is a geodesically convex functional defined on the Wasserstein space, and $\mathcal{F}(\mu)<\infty$ at some $\mu$, then the vector field $\xi=\nabla\frac{\delta\mathcal{F}}{\delta\rho}(\mu)\in L^{2}(\mu;\mathcal{X})$ will satisfy[2]
@@ 3.3 Random update
In the WPCG-R algorithm, we sample the index $j_{l}$ of the coordinate to be updated randomly and uniformly from $[m]$, independently of the previous selections of indexes. To make the convergence rate comparable to the other two schemes, we consider one iteration of WPCG-R as the process of updating $M$ randomly selected coordinates, where the batch size $M$ is of the order $O(m\log m)$. This ensures that with high probability, each coordinate has been updated at least once per iteration. Similar to WPCG-S, WPCG-R is also a descent algorithm regardless of the choice of the step size $\tau$. Pseudo-code of WPCG-R is shown in Algorithm 3.
@@ 3.4.2 Function approximation method
We highlight the optimization problem (15) is unconstrained. By Brenier’s Theorem, the optimal transport map from $\rho^{k}$ to $\rho^{k+1}$ is the gradient of a convex function. Mokrov et al. [35] consider the constrained optimization problem by restricting $T=\nabla\psi$ to the gradient of a convex function $\psi$ and finding the optimal convex function $\psi^{k+1}$ by using input-convex neural networks (ICNN) [3]. On the contrary, Proposition 3.1 shows that an unconstrained optimization problem is enough; directly minimizing the objective functional (15) yields the optimal transport map $T^{k+1}$ from $\rho^{k}$ to $\rho^{k+1}$. The intuition is that, when $T^{k+1}$ is not the optimal transport map, changing $T^{k+1}$ to the optimal transport map $\widetilde{T}^{k+1}$ from $\rho^{k}$ to $\rho^{k+1}$ does not change the first three functional values in (15) but decreases the last term, which contradicts to the optimality of $T^{k+1}$.
In practice, the optimization problem in Proposition 3.1 over the function space can be solved by function approximation methods such as kernel methods and neural networks. Moreover, since we do not have access to the time-evolving density $\rho^{k}$, we still need to approximate the objective functional in Proposition 3.1. Suppose $\{X_{b}^{k}:b\in[B]\}$ are samples drawn from $\rho^{k}$. We provide two concrete examples, which will be used later in Section 5, to show how to approximate the objective functional in Proposition 3.1. The details will be postponed to Appendix C.7.
Example 3.2 (Approximating negative self-entropy).
@@
Example 3.3 (Approximating porous medium type diffusion).
@@
Remark 3.4 (Comparison between FA and SDE approaches).
Compared with the SDE approach, the FA approach can be applied when the diffusion term $\mathcal{H}_{j}$ is beyond the negative self-entropy. According to our numerical results in Section 5, the performance of the FA approach will not be affected when the system contains super-quadratic drift terms. However, smaller step size has to be chosen to apply the SDE approach, since $\tau$ needs to be smaller than $O(L_{\rm c}^{-1})$, where $L_{\rm c}=\max_{j\in[m],x\in\mathcal{X}}|\!|\!|\nabla_{j}^{2}V(x)|\!|\!|_{\rm op}$) is defined in the discussion after Assumption B (Section 4), so that the explicit discretization of the SDE converges. It then takes more iterations for the SDE approach to converge due to this requirement on a smaller step size. An additional attractive aspect of the FA approach is its unbiasedness for any step size $\tau$. This is due to the fact that any fixed-point solution (from a distributional perspective) to its iterative formula must correspond to a global minimum of $\mathcal{F}$. In comparison, the SDE approach often exhibits bias — its fixed point usually deviates by $O(\sqrt{\tau})$ from a global minimum of $\mathcal{F}$. Although choosing a decreasing step size sequence when discretizing the SDE (14) can alleviate the constant bias, it does so at the expense of convergence speed. This is evidenced by the slow convergence observed in Figure 7a, which persists even for the largest permissible step size due to the existence of a super-quadratic error term. ∎
@@ 4 Convergence analysis of WPCG algorithms
The first condition on $h_{j}$ implies that the negative part $h_{j}(\rho_{j})$ is integrable. Details are referred to Remark 3.9 in [1]. The second condition implies the convexity of $h_{j}$ along geodesics [1, 42]. The third condition guarantees that the first variation of $\mathcal{H}_{j}$ at $\rho_{j}$ admits the explicit form as $h_{j}^{\prime}(\rho_{j})$ (Lemma B.1 in Appendix). For example, $h_{j}(x_{j})=x_{j}\log x_{j}$ (corresponding to negative self-entropy) and $h_{j}(x_{j})=x_{j}^{m_{j}}$ with $m_{j}>1$ (corresponding to porous medium type diffusion) satisfy all these three conditions.
@@
This Lipschitz constant is different from the coordinate Lipschitz constant $L_{\rm c}$ used by [50] in $\|\nabla_{j}{V}(y_{j},x_{-j})-\nabla_{j}{V}(z_{j},x_{-j})\|\leq L_{\rm c}\|y_{j}-z_{j}\|$ for all $j\in[m]$. A finite $L_{\rm c}$ guarantees the smoothness of $V$ on each coordinate, which helps control the change of the function value of $V$ after each iteration through the gradient $\nabla V$. [50] also defined the restricted Lipschitz constant $L_{\rm r}$ as $\|\nabla{V}(y_{j},x_{-j})-\nabla{V}(z_{j},x_{-j})\|\leq L_{\rm r}\|y_{j}-z_{j}\|$ to study the convergence property of asynchronous coordinate GD. We also let $L_{\rm g}$ be the global Lipschitz constant defined through $\|\nabla{V}(x)-\nabla{V}(y)\|\leq L_{\rm g}\|x-y\|$, which is commonly used in literature of GD. In fact, we have $0<L/L_{\rm c}\leq\sqrt{m-1}$, $0\leq L/L_{\rm r}\leq\sqrt{1-1/m}$, and $0\leq L/L_{\rm g}\leq 1$. The proof of these connections is provided in Appendix E.
@@
Target functional $\mathcal{F}$ satisfies the following $\lambda$-quadratic growth ($\lambda$-QG) condition: for $\lambda\geq 0$,
@@
It is straightforward to verify that $\mathcal{F}_{j}$ satisfies the $\lambda$-QG condition due to the growth rate of $V_{j}$ (see p.280 in [48] for more details) and is convex along geodesics, but is not strongly convex along geodesics since $V$ is not strongly convex in $[-1,1]^{m}$ (e.g. consider two absolutely continuous measures supported on $[-1,1]^{m}$).
The PL inequality has an analogy on the Wasserstein space which is called log-Sobolev inequality (LSI) [48]. However, it is not easy to directly prove an LSI, mainly due to the following two reasons. Firstly, LSI is often exclusively used to study the KL divergence type objective functionals associated to the relative entropy functional, while we are considering more general objective functionals. Secondly, when we consider a multivariate functional on the Wasserstein space, the multivariate Wasserstein gradient does not tensorize, i.e., the sum of the norm of its blockwise Wasserstein gradient does not equal to the norm of the Wasserstein gradient of the functional treated as a univariate functional (see ahead (18) in Remark 4.10 below).
@@ 4.1 Convergence rates of WPCG under $\lambda$-QG assumption
Theorem 4.2 (Exponential convergence rate of WPCG-P under $\lambda$-QG).
@@
Remark 4.3 (Comments on the step size and iteration complexity of WPCG-P).
@@
i.e., WPCG-P requires at most $O(\frac{\sqrt{m}L}{\lambda}\log({1\over\lambda\varepsilon}))$ iterations to achieve $\varepsilon$-accuracy when the step size $\tau$ is on the scale $O(\frac{1}{L\sqrt{m}})$. [In the sequel, we will drop the logarithmic factor in the iteration complexity when exponential convergence is achieved.] (ii) Both the upper bound of the step size $\tau$ and the smoothness condition of $V$ are necessary for the convergence of the parallel update scheme. The following Example 4.4 shows that the upper bound $O(\frac{1}{L\sqrt{m}})$ of the step size $\tau$ cannot be relaxed, i.e., there exists $V$, $\mathcal{H}_{j}$ and $\mathcal{W}_{j}$ such that WPCG-P diverges when $\tau$ exceeds the scale $O(\frac{1}{L\sqrt{m}})$. Moreover, when we choose $\tau$ on this scale, the iteration complexity is exactly $O(\frac{\sqrt{m}L}{\lambda})$ which coincides with the result in the previous discussion. ∎
Example 4.4 (Optimality of step size and convergence rate of WPCG-P).
@@
Theorem 4.5 (Exponential convergence rate of WPCG-S under $\lambda$-QG).
@@
Remark 4.6 (Comparison with existing results in the Euclidean case).
Similar to the parallel update scheme, by taking $\tau^{-1}=L\sqrt{m-1}$ the iteration complexity also grows at a rate of $O(\frac{L\sqrt{m}}{\lambda})$. When degenerated to optimization on the Euclidean space (by taking $\mathcal{H}_{j}\equiv 0$ and $\mathcal{W}_{j}\equiv 0$), our result is comparable to the rate $O(\frac{\sqrt{m}L_{\rm g}}{\lambda})$ in Theorem 6.3 of [51] for sequential coordinate gradient descent method with strongly convex and smooth functions. When we take $\tau=L^{-1}$, the iteration complexity turns out to be $O(\frac{Lm}{\lambda})$. This iteration complexity is no larger than $O(\frac{\max_{j}L_{j}}{\min_{j}\{L_{j}+\mu_{j}\}}\cdot\frac{mL_{\rm g}}{\lambda})$ derived by Li et al. [33] when applying coordinate proximal gradient descent with the sequential update scheme, where $\lambda_{j}$ and $L_{j}$ are the parameters of strong convexity and smoothness of the $j$-th block, and the step size for updating the $j$-th block is $L_{j}^{-1}$. Their result has an additional factor $\frac{\max_{j}L_{j}}{\min_{j}\{L_{j}+\mu_{j}\}}$ because they approximated the smooth part of the objective function by its first-order Taylor expansion before applying the proximal descent step. As we mentioned, we believe that the difference between the Lipschitz constant in our result and the ones chosen in [51, 33] for optimization problems on the Euclidean space is due to our choice of a proximal-type optimization method. In fact, for the sequential update scheme, the Lipschitz constant can be further relaxed to the lower triangular Lipschitz constant [25]. ∎
Remark 4.7 (Effect of step size in WPCG-S).
The iteration complexity $O(\frac{L\sqrt{m}}{\lambda})$ when $\tau^{-1}={L\sqrt{m-1}}$ in WPCG-S is smaller than the iteration complexity $O(\frac{mL_{\rm g}}{\lambda})$ derived by [44] for alternating minimization algorithm (corresponding to $\tau=\infty$) with the sequential update scheme on the Euclidean space, which shows that using an overly large $\tau$ can still drag down the convergence speed. It is an interesting open problem that if the iteration complexity $O(\frac{L\sqrt{m}}{\lambda})$ can be improved. Even when it degenerates to an optimization problem on $\mathbb{R}^{m}$ ($\mathcal{H}_{j}=\mathcal{W}_{j}=0$ and $\rho^{0}$ is a point mass measure), the optimality of this rate remains as an open problem. [44] showed that there exists an optimization problem (related to the Example 4.4 by choosing $\alpha$ carefully) on $\mathbb{R}^{m}$ such that alternating minimization algorithm with the sequential update scheme has iteration complexity $O(\frac{mL_{\rm g}}{\lambda})$. ∎
@@
Theorem 4.8 (Exponential convergence rate of WPCG-R under $\lambda$-QG).
@@
Remark 4.9 (Comments on the batch size in WPCG-R).
In the proof of Theorem 4.8, we may control the norm of the blockwise Wasserstein gradient by the distance between two consecutive iterates only when all coordinates have been updated at least once in each iteration. For this reason, we need to choose a larger batch size, $M=\lceil 2m\log(mL)\rceil$, compared to the other two update schemes (parallel and sequential), where precisely $M=m$ coordinates are updated in each iteration. ∎
Remark 4.10 (Comparison with existing results in the Euclidean case).
In optimization problems on the Euclidean space, the random update scheme usually has smaller iteration complexity compared with the other two schemes in worst case scenario, such as in coordinate GD [50] and alternating minimization algorithm [44]. However, on the Wasserstein space, we are only able to derive the same iteration complexity up to a log factor as in the other two update schemes, given $\lambda>0$ and a fixed step size $\tau>0$. In particular, when $\tau^{-1}=L\sqrt{2em\log(mL)}$, our result implies an $O(\frac{L\sqrt{m\log(mL)}}{\lambda})$ iteration complexity of WPCG-R, where in each iteration we need to solve $\lceil 2m\log(mL)\rceil$ sub-problems. Recall that there are only $m$ sub-problems in each iteration in WPCG-P and WPCG-S. Therefore, the iteration complexity in WPCG-R is $[\log(mL)]^{3/2}$ times larger than WPCG-P and WPCG-S. This logarithmic factor appears since we need to solve more sub-problems in each iteration compared with the other two update schemes, so that every coordinate has been updated with high probability in each iteration. When $\tau=L^{-1}$, we derive the iteration complexity $O(\frac{mL\log(mL)}{\lambda})$, which is slower than $O(\frac{L_{\rm c}}{\lambda})$ in the coordinate GD with random update scheme derived by Wright and Recht [51] if $mL\log(mL)>L_{\rm c}$. We conjecture that such a sub-optimal rate when specialized to the Euclidean case [44, 51] is due to the extra complexity of dealing with the Wasserstein space, where there is no tensorization structure of the Wasserstein gradient. More precisely, the following key identity

### v2

@@ 2.2 Convex Functionals on Wasserstein Space
only requires the $\lambda$-strong convexity to guarantee its exponential convergence rate (Beck, 2017). Note that a proximal gradient algorithm is also called an implicit gradient descent algorithm, since the first-order optimality condition (FOC) for (8) reads $x^{k+1}_{\text{prox}}=x^{k}_{\text{prox}}-\tau\nabla f(x^{k+1}_{\text{prox}})$. The theoretical advantage of the proximal scheme comes at the price of the additional computational cost of numerically solving this FOC equation.
@@
where $\{\mu_{t}:0\leq t\leq 1\}$ is the constant speed geodesic curve connecting $\mu_{0}$ and $\mu_{1}$. When $\lambda=0$, we simply say that $\mathcal{F}$ is geodesically convex. It can be shown that both functionals $\mathcal{W}_{j}(\rho_{j})$ and $\mathcal{H}_{j}(\rho_{j})$ in our problem formulation (4) are geodesically convex when their defining functions $W_{j}$ and $h_{i}$ are convex (Ambrosio et al., 2005, Chapter 9.3). However, different from the Euclidean case where all convex functions are continuous, geodesically convex functionals may not be continuous with respect to the $\operatorname{W}_{2}$ distance.
Example 3 (Geodesically convex functional may not be continuous).
@@
with the initial distribution $\rho^{0}$ and the step size $\tau>0$. This is analogous to the proximal gradient algorithm (8), i.e., the backward time-discretization of a gradient flow on the Euclidean space. The convergence of Wasserstein proximal gradient scheme for the KL divergence functional $\mathcal{F}_{\operatorname{KL}}(\rho)=\int V\,{\rm d}\rho+\int\rho\log\rho$ is well studied in literature. As on the Euclidean space, piecewise interpolation $\rho_{t}=\rho^{\lfloor\frac{t}{\tau}\rfloor}$ converges to a continuous dynamics on Wasserstein space, i.e., the solution of Fokker–Planck equation $\partial_{t}\rho_{t}=\nabla\cdot(\rho_{t}\nabla(V+\log\rho_{t}))$ (Santambrogio, 2015). Since the Fokker–Planck equation is also the density evolution equation of Langevin dynamics (Jordan et al., 1998)
@@
one can also use the Euler–Maruyama method to numerically approximate the Wasserstein proximal gradient scheme for the KL using particle approximation. It is shown in the literature that when $V$ is strongly convex, the solution $\{\rho_{t}:\,t\geq 0\}$ of the Fokker–Planck equation converges to the unique global minimizer $\rho^{\ast}\,\propto\,e^{-V}$ of $\mathcal{F}_{\rm{KL}}$ (Villani, 2021). Corollary 2.8 in (Yao and Yang, 2022) further shows that the time-discretization of the continuous time dynamic via Wasserstein proximal gradient scheme $\{\rho^{k}:k\in\mathbb{Z}_{+}\}$ is unbiased, and also converges to $\rho^{\ast}\propto e^{-V}$ exponentially fast when $V$ is strongly convex.
@@ 2.3 First Variation and Wasserstein Gradient
for any perturbation $\chi=\tilde{\mu}-\mu$ such that $\tilde{\mu}\in\mathscr{P}_{2}^{r}(\mathcal{X})$ and $\int\,{\rm d}\chi=0$. Clearly, $\frac{\delta\mathcal{F}}{\delta\rho}(\mu)$ is uniquely defined up to additive constant. This is analogous to the gradient of $\mathcal{F}$ in the $L^{2}(\mathcal{X})$ sense. If $\mathcal{F}$ is a geodesically convex functional defined on the Wasserstein space, and $\mathcal{F}(\mu)<\infty$ at some $\mu$, then the vector field $\xi=\nabla\frac{\delta\mathcal{F}}{\delta\rho}(\mu)\in L^{2}(\mu;\mathcal{X})$ will satisfy (Ambrosio et al., 2005)
@@ 3.3 Random Update
In the WPCG-R algorithm, we sample the index $j_{l}$ of the coordinate to be updated randomly and uniformly from $[m]$, independently of the previous selections of indices. To make the convergence rate comparable to the other two schemes, we consider one iteration of WPCG-R as the process of updating $M$ randomly selected coordinates, where the batch size $M$ is of the order $O(m\log m)$. This ensures that with high probability, each coordinate has been updated at least once per iteration. Similar to WPCG-S, WPCG-R is also a descent algorithm regardless of the choice of the step size $\tau$. Pseudo-code of WPCG-R is shown in Algorithm 3.
@@ 3.4.2 Function Approximation Method
We highlight the optimization problem (15) is unconstrained. By Brenier’s Theorem, the optimal transport map from $\rho^{k}$ to $\rho^{k+1}$ is the gradient of a convex function. Mokrov et al. (2021) consider the constrained optimization problem by restricting $T=\nabla\psi$ to the gradient of a convex function $\psi$ and finding the optimal convex function $\psi^{k+1}$ by using input-convex neural networks (ICNN) (Amos et al., 2017). On the contrary, Proposition 4 shows that an unconstrained optimization problem is enough; directly minimizing the objective functional (15) yields the optimal transport map $T^{k+1}$ from $\rho^{k}$ to $\rho^{k+1}$. The intuition is that, when $T^{k+1}$ is not the optimal transport map, changing $T^{k+1}$ to the optimal transport map $\widetilde{T}^{k+1}$ from $\rho^{k}$ to $\rho^{k+1}$ does not change the first three functional values in (15) but decreases the last term, which contradicts to the optimality of $T^{k+1}$.
In practice, the optimization problem in Proposition 4 over the function space can be solved by function approximation methods such as kernel methods and neural networks. Moreover, since we do not have access to the time-evolving density $\rho^{k}$, we still need to approximate the objective functional in Proposition 4. Suppose $\{X_{b}^{k}:b\in[B]\}$ are samples drawn from $\rho^{k}$. We provide two concrete examples, which will be used later in Section 5, to show how to approximate the objective functional in Proposition 4. The details will be postponed to Appendix C.7.
Example 5 (Approximating negative self-entropy).
@@
Example 6 (Approximating porous medium type diffusion).
@@
Remark 7 (Comparison between FA and SDE approaches).
Compared with the SDE approach, the FA approach can be applied when the diffusion term $\mathcal{H}_{j}$ is beyond the negative self-entropy. According to our numerical results in Section 5, the performance of the FA approach will not be affected when the system contains super-quadratic drift terms. However, smaller step size has to be chosen to apply the SDE approach, since $\tau$ needs to be smaller than $O(L_{\rm c}^{-1})$, where $L_{\rm c}=\max_{j\in[m],x\in\mathcal{X}}|\!|\!|\nabla_{j}^{2}V(x)|\!|\!|_{\rm op}$) is defined in the discussion after Assumption B (Section 4), so that the explicit discretization of the SDE converges. It then takes more iterations for the SDE approach to converge due to this requirement on a smaller step size. An additional attractive aspect of the FA approach is its unbiasedness for any step size $\tau$. This is due to the fact that any fixed-point solution (from a distributional perspective) to its iterative formula must correspond to a global minimum of $\mathcal{F}$. In comparison, the SDE approach often exhibits bias—its fixed point usually deviates by $O(\sqrt{\tau})$ from a global minimum of $\mathcal{F}$. Although choosing a decreasing step size sequence when discretizing the SDE (14) can alleviate the constant bias, it does so at the expense of convergence speed. This is evidenced by the slow convergence observed in Figure 7(a), which persists even for the largest permissible step size due to the existence of a super-quadratic error term. ∎
@@ 4 Convergence Analysis of WPCG Algorithms
The first condition on $h_{j}$ implies that the negative part $h_{j}(\rho_{j})$ is integrable. Details are referred to Remark 3.9 in (Ambrosio and Savaré, 2007). The second condition implies the convexity of $h_{j}$ along geodesics (Ambrosio and Savaré, 2007; Santambrogio, 2015). The third condition guarantees that the first variation of $\mathcal{H}_{j}$ at $\rho_{j}$ admits the explicit form as $h_{j}^{\prime}(\rho_{j})$ (Lemma 9 in Appendix). For example, $h_{j}(x_{j})=x_{j}\log x_{j}$ (corresponding to negative self-entropy) and $h_{j}(x_{j})=x_{j}^{m_{j}}$ with $m_{j}>1$ (corresponding to porous medium type diffusion) satisfy all these three conditions.
@@
This Lipschitz constant is different from the coordinate Lipschitz constant $L_{\rm c}$ used by Wright (2015) in $\|\nabla_{j}{V}(y_{j},x_{-j})-\nabla_{j}{V}(z_{j},x_{-j})\|\leq L_{\rm c}\|y_{j}-z_{j}\|$ for all $j\in[m]$. A finite $L_{\rm c}$ guarantees the smoothness of $V$ on each coordinate, which helps control the change of the function value of $V$ after each iteration through the gradient $\nabla V$. Wright (2015) also defined the restricted Lipschitz constant $L_{\rm r}$ as $\|\nabla{V}(y_{j},x_{-j})-\nabla{V}(z_{j},x_{-j})\|\leq L_{\rm r}\|y_{j}-z_{j}\|$ to study the convergence property of asynchronous coordinate GD. We also let $L_{\rm g}$ be the global Lipschitz constant defined through $\|\nabla{V}(x)-\nabla{V}(y)\|\leq L_{\rm g}\|x-y\|$, which is commonly used in literature of GD. In fact, we have $0<L/L_{\rm c}\leq\sqrt{m-1}$, $0\leq L/L_{\rm r}\leq\sqrt{1-1/m}$, and $0\leq L/L_{\rm g}\leq 1$. The proof of these connections is provided in Appendix E.
@@
The target functional $\mathcal{F}$ satisfies the following $\lambda$-quadratic growth ($\lambda$-QG) condition: for $\lambda\geq 0$,
@@
It is straightforward to verify that $\mathcal{F}_{j}$ satisfies the $\lambda$-QG condition due to the growth rate of $V_{j}$ (see p.280 in (Villani, 2021) for more details) and is convex along geodesics, but is not strongly convex along geodesics since $V$ is not strongly convex in $[-1,1]^{m}$ (e.g. consider two absolutely continuous measures supported on $[-1,1]^{m}$).
The PL inequality has an analogy on the Wasserstein space which is called log-Sobolev inequality (LSI) (Villani, 2021). However, it is not easy to directly prove an LSI, mainly due to the following two reasons. Firstly, LSI is often exclusively used to study the KL divergence type objective functionals associated to the relative entropy functional, while we are considering more general objective functionals. Secondly, when we consider a multivariate functional on the Wasserstein space, the multivariate Wasserstein gradient does not tensorize, i.e., the sum of the norm of its blockwise Wasserstein gradient does not equal to the norm of the Wasserstein gradient of the functional treated as a univariate functional (see ahead (18) in Remark 17 below).
@@ 4.1 Convergence Rates of WPCG under $\lambda$-QG Assumption
Theorem 9 (Exponential convergence rate of WPCG-P under $\lambda$-QG).
@@
Remark 10 (Comments on the step size and iteration complexity of WPCG-P).
@@
i.e., WPCG-P requires at most $O(\frac{\sqrt{m}L}{\lambda}\log({1\over\lambda\varepsilon}))$ iterations to achieve $\varepsilon$-accuracy when the step size $\tau$ is on the scale $O(\frac{1}{L\sqrt{m}})$. [In the sequel, we will drop the logarithmic factor in the iteration complexity when exponential convergence is achieved.] (ii) Both the upper bound of the step size $\tau$ and the smoothness condition of $V$ are necessary for the convergence of the parallel update scheme. The following Example 11 shows that the upper bound $O(\frac{1}{L\sqrt{m}})$ of the step size $\tau$ cannot be relaxed, i.e., there exists $V$, $\mathcal{H}_{j}$, and $\mathcal{W}_{j}$ such that WPCG-P diverges when $\tau$ exceeds the scale $O(\frac{1}{L\sqrt{m}})$. Moreover, when we choose $\tau$ on this scale, the iteration complexity is exactly $O(\frac{\sqrt{m}L}{\lambda})$ which coincides with the result in the previous discussion. ∎
Example 11 (Optimality of step size and convergence rate of WPCG-P).
@@
Theorem 12 (Exponential convergence rate of WPCG-S under $\lambda$-QG).
@@
Remark 13 (Comparison with existing results in the Euclidean case).
Similar to the parallel update scheme, by taking $\tau^{-1}=L\sqrt{m-1}$ the iteration complexity also grows at a rate of $O(\frac{L\sqrt{m}}{\lambda})$. When degenerated to optimization on the Euclidean space (by taking $\mathcal{H}_{j}\equiv 0$ and $\mathcal{W}_{j}\equiv 0$), our result is comparable to the rate $O(\frac{\sqrt{m}L_{\rm g}}{\lambda})$ in Theorem 6.3 of (Wright and Recht, 2022) for sequential coordinate gradient descent method with strongly convex and smooth functions. When we take $\tau=L^{-1}$, the iteration complexity turns out to be $O(\frac{Lm}{\lambda})$. This iteration complexity is no larger than $O(\frac{\max_{j}L_{j}}{\min_{j}\{L_{j}+\mu_{j}\}}\cdot\frac{mL_{\rm g}}{\lambda})$ derived by Li et al. (2017) when applying coordinate proximal gradient descent with the sequential update scheme, where $\lambda_{j}$ and $L_{j}$ are the parameters of strong convexity and smoothness of the $j$-th block, and the step size for updating the $j$-th block is $L_{j}^{-1}$. Their result has an additional factor $\frac{\max_{j}L_{j}}{\min_{j}\{L_{j}+\mu_{j}\}}$ because they approximated the smooth part of the objective function by its first-order Taylor expansion before applying the proximal descent step. As we mentioned, we believe that the difference between the Lipschitz constant in our result and the ones chosen in (Wright and Recht, 2022; Li et al., 2017) for optimization problems on the Euclidean space is due to our choice of a proximal-type optimization method. In fact, for the sequential update scheme, the Lipschitz constant can be further relaxed to the lower triangular Lipschitz constant (Hua and Yamashita, 2015). ∎
Remark 14 (Effect of step size in WPCG-S).
The iteration complexity $O(\frac{L\sqrt{m}}{\lambda})$ when $\tau^{-1}={L\sqrt{m-1}}$ in WPCG-S is smaller than the iteration complexity $O(\frac{mL_{\rm g}}{\lambda})$ derived by Sun and Ye (2021) for alternating minimization algorithm (corresponding to $\tau=\infty$) with the sequential update scheme on the Euclidean space, which shows that using an overly large $\tau$ can still drag down the convergence speed. It is an interesting open problem that if the iteration complexity $O(\frac{L\sqrt{m}}{\lambda})$ can be improved. Even when it degenerates to an optimization problem on $\mathbb{R}^{m}$ ($\mathcal{H}_{j}=\mathcal{W}_{j}=0$ and $\rho^{0}$ is a point mass measure), the optimality of this rate remains as an open problem. Sun and Ye (2021) showed that there exists an optimization problem (related to the Example 11 by choosing $\alpha$ carefully) on $\mathbb{R}^{m}$ such that alternating minimization algorithm with the sequential update scheme has iteration complexity $O(\frac{mL_{\rm g}}{\lambda})$. ∎
@@
Theorem 15 (Exponential convergence rate of WPCG-R under $\lambda$-QG).
@@
Remark 16 (Comments on the batch size in WPCG-R).
In the proof of Theorem 15, we may control the norm of the blockwise Wasserstein gradient by the distance between two consecutive iterates only when all coordinates have been updated at least once in each iteration. For this reason, we need to choose a larger batch size, $M=\lceil 2m\log(mL)\rceil$, compared to the other two update schemes (parallel and sequential), where precisely $M=m$ coordinates are updated in each iteration. ∎
Remark 17 (Comparison with existing results in the Euclidean case).
In optimization problems on the Euclidean space, the random update scheme usually has smaller iteration complexity compared with the other two schemes in worst case scenario, such as in coordinate GD (Wright, 2015) and alternating minimization algorithm (Sun and Ye, 2021). However, on the Wasserstein space, we are only able to derive the same iteration complexity up to a log factor as in the other two update schemes, given $\lambda>0$ and a fixed step size $\tau>0$. In particular, when $\tau^{-1}=L\sqrt{2em\log(mL)}$, our result implies an $O(\frac{L\sqrt{m\log(mL)}}{\lambda})$ iteration complexity of WPCG-R, where in each iteration we need to solve $\lceil 2m\log(mL)\rceil$ sub-problems. Recall that there are only $m$ sub-problems in each iteration in WPCG-P and WPCG-S. Therefore, the iteration complexity in WPCG-R is $[\log(mL)]^{3/2}$ times larger than WPCG-P and WPCG-S. This logarithmic factor appears since we need to solve more sub-problems in each iteration compared with the other two update schemes, so that every coordinate has been updated with high probability in each iteration. When $\tau=L^{-1}$, we derive the iteration complexity $O(\frac{mL\log(mL)}{\lambda})$, which is slower than $O(\frac{L_{\rm c}}{\lambda})$ in the coordinate GD with random update scheme derived by Wright and Recht (2022) if $mL\log(mL)>L_{\rm c}$. We conjecture that such a sub-optimal rate when specialized to the Euclidean case (Sun and Ye, 2021; Wright and Recht, 2022) is due to the extra complexity of dealing with the Wasserstein space, where there is no tensorization structure of the Wasserstein gradient. More precisely, the following key identity